OpenStudy (marcon):

Given the function f(x) = 3(x – 3)^2 + 2, indicate the shifts that will affect the location of the vertex, and explain what effect they will have. Use complete sentences. f(x+4) f(x) + 4 f(4x) 4⋅f(x)

OpenStudy (marcon):

@Nurali @phi

OpenStudy (marcon):

@iGreen

OpenStudy (anonymous):

y=f(x)

OpenStudy (anonymous):

e f(x)=x

OpenStudy (anonymous):

ignore the e on the second answer

OpenStudy (anonymous):

, a straight line going through the origin and up and to the right with a slope of 1.

OpenStudy (anonymous):

Now let's add 1 to the function: y=f(x)+1=x+1

OpenStudy (marcon):

Why would we add a 1?

OpenStudy (anonymous):

we're changing the function and the graph

OpenStudy (anonymous):

Adding a positive value to the result translates the graph in the direction of positive y, and adding a negative value translates in the direction of negative y.

OpenStudy (marcon):

Oh okay. Are we doing the first one? f(x+4) ?

OpenStudy (anonymous):

So we're trying to determine if we shift from left or right

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

What is the value of y at x=0? x=1, x=−3?

OpenStudy (marcon):

We shift to the left 4 units. Correct?

OpenStudy (anonymous):

Yes! :)

OpenStudy (marcon):

So the first one would be (-1,2) right?

OpenStudy (anonymous):

You are indeed correct

OpenStudy (marcon):

Alright I'm kinda getting this now. For f(x)+4 would be move 4 units horizontally?

OpenStudy (anonymous):

Very good :)

OpenStudy (marcon):

Idk how to do f(4x)...lol

OpenStudy (anonymous):

ok, let me help

OpenStudy (anonymous):

we have y=f(x)=x2. y=f(x+3)+2=(x+3)2+2 and y=f(x+2)−3=(x+2)2−3

OpenStudy (anonymous):

y=f(x)+(−1)=f(x)−1

OpenStudy (anonymous):

y=f(x)+a translates the graph along the y axis by a, with a>0 moving the graph in the direction of increasing y values, and a<0 moving the graph in the direction of decreasing y values. What happens if a=0?

OpenStudy (marcon):

Idk... this is kinda confusing now

OpenStudy (anonymous):

ok, I'm going to call extra help

OpenStudy (anonymous):

If you don't mind

OpenStudy (marcon):

I think we would shift the function to (1,2)

OpenStudy (marcon):

or (3/4 , 2) <---- I think this one is correct

OpenStudy (anonymous):

Yes, that is correct

OpenStudy (marcon):

We transform the function to g(x), which would then shift the function to (3/4 , 2)... Something like that

OpenStudy (marcon):

What about the last one? My guess is we also transform the function to g(x) like from the third one, then shift the function to (3,8)

OpenStudy (anonymous):

Yes

OpenStudy (marcon):

Is that it?

OpenStudy (anonymous):

yes

OpenStudy (marcon):

Would you mind helping me with a few more?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

just repost the question and tag me

OpenStudy (marcon):

alright

OpenStudy (marcon):

I posted it and tagged you